OpenStudy (anonymous):

A giant watermelon weighed 100 pounds and was 99% water by weight in pounds. While standing in the sun, some water evaporated, so that the watermelon was only 98% water by weight in pounds. How many pounds did the watermelon then weigh?

OpenStudy (anonymous):

hmm

OpenStudy (foolaroundmath):

Initially: Weight of non-water substance = 1 pound (1% of 100 pounds) Weight of water = 99 pounds Suppose 'x' pounds of water evaporated. Now, weight of water = 99 -x pounds Weight of non-water substance remains same as before = 1 pound Percentage of water now = weight of water/total weight = (99-x)*100/(99-x+1) This is given to be 98 Thus, \[\frac{99-x}{100-x} = \frac{98}{100}\] see if you can solve this

OpenStudy (anonymous):

9900 - 100x = 9800 - 98x 100 = 2x x = 50?

OpenStudy (foolaroundmath):

yes .. x is the weight of water evaporated, new weight = 99-x +1 = 100-x = 50 pounds

OpenStudy (anonymous):

Alright. Thank you!